OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = -3/8 and passing through (5, -4)

OpenStudy (saifoo.khan):

c = -17/8

OpenStudy (saifoo.khan):

y = mx + b y = -3x/8 -17/8 8y = -3x -17